Kevin's income elasticity for good A is equal to -1.5. His current income is $40,000 per year and he buys 200 units of good A annually. If his income falls to $36,000 how many units of good A will he purchase?
Answer #1

% decrease in income = (36000-40000)/40000 = -10%

Income elasticity of demand = % change in quantity demanded / % change in income

-1.5 = % change in quantity demanded/(-10%)

% change in quantity demanded = 15%

So,

Units of good A to be purchased = 200*(1+15%)

Units of good A to be purchased = 230 units